# Neural Dueling Bandits

### v1

@@ 7 Conclusion
Due to their prevalence in many real-life applications, from online recommendations to ranking web search results, we consider contextual dueling bandit problems that can have a complex and non-linear reward function. We used a neural network to estimate this reward function using preference feedback observed for the previously selected arms. We proposed upper confidence bound- and Thompson sampling-based algorithms with sub-linear regret guarantees for contextual dueling bandits. Experimental results using synthetic functions corroborate our theoretical results. Our algorithms and theoretical results can also provide insights into the celebrated reinforcement learning with human feedback (RLHF) algorithm, such as a theoretical guarantee on the quality of the learned reward model. We also extend our results to contextual bandit problems with binary feedback, which is in itself a non-trivial contribution. A limitation of our work is that we currently do not account for problems where multiple arms are selected simultaneously (multi-way preference), which is an interesting future direction. Another future topic is to apply our algorithms to important real-world problems involving preference or binary feedback, e.g., LLM alignment using human feedback.
@@ Appendix A Theoretical Analysis for NDB-UCB
Here, we first list the specific conditions we need for the width $m$ of the NN:
@@ A.1 Theoretical Guarantee about the Neural Network
Now, 2 allows us to obtain the following lemmas regarding the gradients of the NN.
@@
It can be easily verified that our $\tau=2\sqrt{\frac{t}{m\lambda}}$ satisfies the requirement on $\tau$ specified in Lemmas B.5 and B.6 from [8]. Therefore, the results from Lemmas B.5 and B.6 from [8] are applicable for $\theta_{t}$ because our 2 guarantees that $\left\|\theta_{t}-\theta_{0}\right\|_{2}\leq\tau$. ∎
In addition, Lemmas B.4 from [7] allows us to obtain the following lemma, which shows that the output of the NN can be approximated by its linearization.
@@
$|h(x;\theta_{t})-\langle\theta_{t}-\theta_{0},g(x;\theta_{0})\rangle|\leq C_{2}\tau^{4/3}L^{3}\sqrt{m\log m}$ $=C_{2}m^{-1/6}\sqrt{\log m}L^{3}\left(\frac{t}{\lambda}\right)^{4/3}$
$=\varepsilon^{\prime}_{m,t},$
@@
An immediate consequence of 4 is the following lemma.
Lemma 5.
For all $t\in[T]$, we have for all $x,x^{\prime}\in\mathcal{X}_{t}$ that
$|\langle\varphi(x)-\varphi(x^{\prime}),\theta_{t}-\theta_{0}\rangle-(h(x;\theta_{t})-h(x^{\prime};\theta_{t}))|\leq 2\varepsilon^{\prime}_{m,t}.$
@@
By re-arranging the left-hand side and then using 4, we get
@@ A.2.1 Proof of 6
Note that we choose $\lambda$ such that $\frac{\lambda}{\kappa_{\mu}}>1$. This allows us to show that $V_{t-1}\succeq\frac{\lambda}{\kappa_{\mu}}I$ and hence $V_{t-1}^{-1}\preceq\frac{\kappa_{\mu}}{\lambda}I$. Recall that 1 tells us that $\left\|\theta_{f}-\theta_{0}\right\|_{2}\leq\frac{B}{\sqrt{m}}$, which tells us that
@@
in which the last inequality follows from 3. Now the norm of $A_{1}$ can be bounded as:
@@
in which we have used 4 in the last inequality. Also, 3 allows us to show that $\left\|\widetilde{\varphi}^{\prime}_{s}\right\|_{2}=\left\|g(x_{s,1};\theta_{t})-g(x_{s,2};\theta_{t})\right\|_{2}\leq\left\|g(x_{s,1};\theta_{t})\right\|_{2}+\left\|g(x_{s,2};\theta_{t})\right\|_{2}\leq 2C_{3}\sqrt{mL}$.
@@
Next, making use of the $1$-sub-sub-Gaussianity of the sequence of noise $\{\varepsilon_{s}\}$ and Theorem 1 from [18], we can show that with probability of at least $1-\delta$,
@@ A.2.2 Proof of 1
Denote $\varphi(x)=g(x;\theta_{0})$. Recall that 1 tells us that $f(x)=\langle g(x;\theta_{0}),\theta_{f}-\theta_{0}\rangle=\langle\varphi(x),\theta_{f}-\theta_{0}\rangle$ for all $x\in\mathcal{X}_{t},t\in[T]$. To begin with, for all $x,x^{\prime}\in\mathcal{X}_{t},t\in[T]$ we have that
@@
in which we have used 6 in the last inequality. Now making use of the equation above and 5, we have that
@@ A.3 Regret Analysis
Step $(a)$ follows from Eq. 26, step $(b)$ results from 5, step $(c)$ follows from the way in which $x_{t,2}$ is selected: $x_{t,2}=\arg\max_{x\in\mathcal{X}_{t}}\left(h(x;\theta_{t})+\left\|\frac{1}{\sqrt{m}}\left(\varphi(x)-\varphi(x_{t,1})\right)\right\|_{V_{t-1}^{-1}}\left(\beta_{T}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}+1\right)\right)$, and step $(d)$ follows from the way in which $x_{t,1}$ is selected: $x_{t,1}=\arg\max_{x\in\mathcal{X}_{t}}h(x;\theta_{t})$.
@@
in which we have denoted $c_{0}>0$ as an absolute constant such that $\frac{1}{m}\left\|\varphi(x)-\varphi(x^{\prime})\right\|_{2}^{2}\leq c_{0},\forall x,x^{\prime}\in\mathcal{X}_{t},t\in[T]$. Note that this is similar to the standard assumption in the literature that the value of the NTK is upper-bounded by a constant [47]. Therefore, this implies that $\sigma_{t-1}^{2}(x_{t,1},x_{t,2})/c_{0}\leq 1$ for some constant $c_{0}\geq 1$. Recall that we choose $\lambda$ such that $\lambda/\kappa_{\mu}\geq 1$. Note that for any $\alpha\in[0,1]$, we have that $\alpha/2\leq\log(1+\alpha)$. With these, we have that
@@
Following the analysis of [14] and using the chain rule of conditional information gain, we can show that
@@
$\mathfrak{R}_{T}$ $\leq\frac{3}{2}\left(\beta_{T}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}+1\right)\sqrt{T2c_{0}\widetilde{d}}+1$
$=\widetilde{\mathcal{O}}\left(\left(\frac{\sqrt{\widetilde{d}}}{\kappa_{\mu}}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{\widetilde{d}T}\right).$
@@ Appendix B Theoretical Analysis for NDB-TS
$\begin{split}f(x_{t,2})-f(x_{t,1})&=h(x_{t,2};\theta_{t})-h(x_{t,1};\theta_{t})+\nu_{T}\sigma_{t-1}(x_{t,2},x_{t,1})+2\varepsilon^{\prime}_{m,t}\\ &\leq\nu_{T}\sigma_{t-1}(x_{t,2},x_{t,1})+2\varepsilon^{\prime}_{m,t}\\ &\leq c_{t}\sigma_{t-1}(x_{t,2},x_{t,1})+2\varepsilon^{\prime}_{m,t},\end{split}$ (34)
in which the first inequality follows from 9, and the second inequality follows because $\nu_{T}\leq c_{t}$ by definition. Now we can plug Eq. 33 and Eq. 34 into Eq. 32:
@@
$=\widetilde{O}\left(\left(\sqrt{\frac{\widetilde{d}}{\kappa_{\mu}\lambda}}+B\right)\sqrt{T\frac{\lambda}{\kappa_{\mu}}\widetilde{d}}\right)$
$=\widetilde{O}\left(\left(\frac{\sqrt{\widetilde{d}}}{\kappa_{\mu}}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{T\widetilde{d}}\right).$
@@
Appendix C Theoretical Analysis for Neural Contextual Bandits with Binary Feedback (Section 5)

### v2

@@ 7 Conclusion
Due to their prevalence in many real-life applications, from online recommendations to ranking web search results, we consider contextual dueling bandit problems that can have a complex and non-linear reward function. We used a neural network to estimate this reward function using human preference feedback observed for the previously selected arms. We proposed upper confidence bound- and Thompson sampling-based algorithms with sub-linear regret guarantees for contextual dueling bandits. Experimental results using synthetic functions corroborate our theoretical results. Our algorithms and theoretical results can also provide insights into the celebrated reinforcement learning with human feedback (RLHF) algorithm, such as a theoretical guarantee on the quality of the learned reward model. We also extend our results to contextual bandit problems with binary feedback, which is in itself a non-trivial contribution. A limitation of our work is that we currently do not account for problems where multiple arms are selected simultaneously (multi-way preference), which is an interesting future direction. Another future topic is to apply our algorithms to important real-world problems involving preference or binary feedback, e.g., LLM alignment using human feedback.
Ethics Statement
The contributions of our work are primarily theoretical. Therefore, we do not foresee any immediate negative societal impact in the short term. Regarding our longer-term impacts, as discussed in Appendix E, our algorithms can be potentially adopted to improve online RLHF. On the positive side, our work can lead to better and more efficient alignment of LLMs through improved online RLHF, which could benefit society. On the other hand, the potential negative societal impacts arising from RLHF may also apply to our work. On the other hand, potential mitigation measures to prevent the misuse of RLHF would also help safeguard the potential misuse of our algorithms.
Acknowledgements
This research is supported by the National Research Foundation (NRF), Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme. The Mens, Manus, and Machina (M3S) is an interdisciplinary research group (IRG) of the Singapore MIT Alliance for Research and Technology (SMART) centre. DesCartes: this research is supported by the National Research Foundation, Prime Minister’s Office, Singapore under its Campus for Research Excellence and Technological Enterprise (CREATE) programme.
@@ Appendix A Theoretical Analysis for NDB-UCB
Let $\{x_{(n)}\}_{n=1}^{TK}$ be a set of all possible context-arm feature vectors: $\{x_{t,a}\}_{1\leq t\leq T,1\leq a\leq K}$, where $n=K(t-1)+a$. Define
$\widetilde{\mathbf{H}}_{p,q}^{(1)}=\mathbf{\Sigma}_{p,q}^{(1)}=\langle x_{(p)},x_{(q)}\rangle,\\ \mathbf{A}_{p,q}^{(l)}=\begin{pmatrix}\mathbf{\Sigma}_{p,q}^{(l)}&\mathbf{\Sigma}_{p,q}^{(l)}&\\ \mathbf{\Sigma}_{p,q}^{(l)}&\mathbf{\Sigma}_{q,q}^{(l)}&\end{pmatrix},$
$\mathbf{\Sigma}_{p,q}^{(l+1)}=2\mathbb{E}_{(u,v)\sim\mathcal{N}(0,\mathbf{A}_{p,q}^{(l)})}[\max\{u,0\}\max\{v,0\}],$
$\widetilde{\mathbf{H}}_{p,q}^{(l+1)}=2\widetilde{\mathbf{H}}_{p,q}^{(l)}\mathbb{E}_{(u,v)\sim\mathcal{N}(0,\mathbf{A}_{p,q}^{(l)})}[\mathbb{1}(u\geq 0)\mathbb{1}(v\geq 0)]+\mathbf{\Sigma}_{p,q}^{(l+1)}.$
Then, $\mathbf{H}=(\widetilde{\mathbf{H}}^{(L)}+\mathbf{\Sigma}^{(L)})/2$ is called the neural tangent kernel (NTK) matrix on the set of context-arm feature vectors $\{x_{(n)}\}_{n=1}^{TK}$.
Next, we first list the specific conditions we need for the width $m$ of the NN:
@@ A.1 Theoretical Guarantee about the Neural Network
Now, 3 allows us to obtain the following lemmas regarding the gradients of the NN.
@@
It can be easily verified that our $\tau=2\sqrt{\frac{t}{m\lambda}}$ satisfies the requirement on $\tau$ specified in Lemmas B.5 and B.6 from Zhang et al. (2021). Therefore, the results from Lemmas B.5 and B.6 from Zhang et al. (2021) are applicable for $\theta_{t}$ because our 3 guarantees that $\left\|\theta_{t}-\theta_{0}\right\|_{2}\leq\tau$. ∎
In addition, Lemmas B.4 from Zhou et al. (2020) allows us to obtain the following lemma, which shows that the output of the NN can be approximated by its linearization.
@@
$|h(x;\theta_{t})-\langle\theta_{t}-\theta_{0},g(x;\theta_{0})\rangle|\leq C_{2}\tau^{4/3}L^{3}\sqrt{m\log m}$ $=C_{2}m^{-1/6}\sqrt{\log m}L^{3}\left(\frac{t}{\lambda}\right)^{4/3}=\varepsilon^{\prime}_{m,t},$
@@
An immediate consequence of 5 is the following lemma. See 1
@@
By re-arranging the left-hand side and then using 5, we get
@@ A.2.1 Proof of 6
Note that we choose $\lambda$ such that $\frac{\lambda}{\kappa_{\mu}}>1$. This allows us to show that $V_{t-1}\succeq\frac{\lambda}{\kappa_{\mu}}I$ and hence $V_{t-1}^{-1}\preceq\frac{\kappa_{\mu}}{\lambda}I$. Recall that 2 tells us that $\left\|\theta_{f}-\theta_{0}\right\|_{2}\leq\frac{B}{\sqrt{m}}$, which tells us that
@@
in which the last inequality follows from 4. Now the norm of $A_{1}$ can be bounded as:
@@
in which we have used 5 in the last inequality. Also, 4 allows us to show that $\left\|\widetilde{\varphi}^{\prime}_{s}\right\|_{2}=\left\|g(x_{s,1};\theta_{t})-g(x_{s,2};\theta_{t})\right\|_{2}\leq\left\|g(x_{s,1};\theta_{t})\right\|_{2}+\left\|g(x_{s,2};\theta_{t})\right\|_{2}\leq 2C_{3}\sqrt{mL}$.
@@
Next, making use of the $1$-sub-sub-Gaussianity of the sequence of noise $\{\varepsilon_{s}\}$ and Theorem 1 from Abbasi-Yadkori et al. (2011), we can show that with probability of at least $1-\delta$,
@@ A.2.2 Proof of 1
Denote $\varphi(x)=g(x;\theta_{0})$. Recall that 2 tells us that $f(x)=\langle g(x;\theta_{0}),\theta_{f}-\theta_{0}\rangle=\langle\varphi(x),\theta_{f}-\theta_{0}\rangle$ for all $x\in\mathcal{X}_{t},t\in[T]$. To begin with, for all $x,x^{\prime}\in\mathcal{X}_{t},t\in[T]$ we have that
@@
in which we have used 6 in the last inequality. Now making use of the equation above and 1, we have that
@@ A.3 Regret Analysis
Step $(a)$ follows from Eq. 24, step $(b)$ results from 1, step $(c)$ follows from the way in which $x_{t,2}$ is selected: $x_{t,2}=\arg\max_{x\in\mathcal{X}_{t}}\left(h(x;\theta_{t})+\left\|\frac{1}{\sqrt{m}}\left(\varphi(x)-\varphi(x_{t,1})\right)\right\|_{V_{t-1}^{-1}}\left(\beta_{T}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}+1\right)\right)$, and step $(d)$ follows from the way in which $x_{t,1}$ is selected: $x_{t,1}=\arg\max_{x\in\mathcal{X}_{t}}h(x;\theta_{t})$.
@@
in which we have denoted $c_{0}>0$ as an absolute constant such that $\frac{1}{m}\left\|\varphi(x)-\varphi(x^{\prime})\right\|_{2}^{2}\leq c_{0},\forall x,x^{\prime}\in\mathcal{X}_{t},t\in[T]$. Note that this is similar to the standard assumption in the literature that the value of the NTK is upper-bounded by a constant (Kassraie & Krause, 2022). Therefore, this implies that $\sigma_{t-1}^{2}(x_{t,1},x_{t,2})/c_{0}\leq 1$ for some constant $c_{0}\geq 1$. Recall that we choose $\lambda$ such that $\lambda/\kappa_{\mu}\geq 1$. Note that for any $\alpha\in[0,1]$, we have that $\alpha/2\leq\log(1+\alpha)$. With these, we have that
@@
Following the analysis of Chowdhury & Gopalan (2017) and using the chain rule of conditional information gain, we can show that
@@
$\mathfrak{R}_{T}$ $\leq\frac{3}{2}\left(\beta_{T}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}+1\right)\sqrt{T2c_{0}\widetilde{d}}+1=\widetilde{\mathcal{O}}\left(\left(\frac{\sqrt{\widetilde{d}}}{\kappa_{\mu}}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{\widetilde{d}T}\right).$
@@ Appendix B Theoretical Analysis for NDB-TS
$\begin{split}f(x_{t,2})-f(x_{t,1})&\leq h(x_{t,2};\theta_{t})-h(x_{t,1};\theta_{t})+\nu_{T}\sigma_{t-1}(x_{t,2},x_{t,1})+2\varepsilon^{\prime}_{m,t}\\ &\leq\nu_{T}\sigma_{t-1}(x_{t,2},x_{t,1})+2\varepsilon^{\prime}_{m,t}\\ &\leq c_{t}\sigma_{t-1}(x_{t,2},x_{t,1})+2\varepsilon^{\prime}_{m,t},\end{split}$ (32)
in which the first inequality follows from 9, the second inequality results from the way in which $x_{t,1}$ is selected: $x_{t,1}=\arg\max_{x\in\mathcal{X}_{t}}h(x;\theta_{t})$, and the third inequality follows because $\nu_{T}\leq c_{t}$ by definition. Now we can plug Eq. 31 and Eq. 32 into Eq. 30:
@@
$=\widetilde{O}\left(\left(\sqrt{\frac{\widetilde{d}}{\kappa_{\mu}\lambda}}+B\right)\sqrt{T\frac{\lambda}{\kappa_{\mu}}\widetilde{d}}\right)=\widetilde{O}\left(\left(\frac{\sqrt{\widetilde{d}}}{\kappa_{\mu}}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{T\widetilde{d}}\right).$
@@
Appendix C Additional experiments for Section 5
(a) $10(x^{\top}\theta)^{2}$ (Average)
(b) $10(x^{\top}\theta)^{2}$ (Weak)
(c) $20(x^{\top}\theta)^{2}$ (Average)
(d) $20(x^{\top}\theta)^{2}$ (Weak)
(e) $30(x^{\top}\theta)^{2}$ (Average)
(f) $30(x^{\top}\theta)^{2}$ (Weak)
(g) $40(x^{\top}\theta)^{2}$ (Average)
(h) $40(x^{\top}\theta)^{2}$ (Weak)
Figure 5: Comparisons of cumulative regret (average and weak) of different dueling bandits algorithms for different form of Square function, i.e., $a(x^{\top}\theta)^{2}$, where $a=\{10,20,30,40\}$.
(a) $cos(3x^{\top}\theta)$ (Average)
(b) $cos(3x^{\top}\theta)$ (Weak)
(c) $3cos(x^{\top}\theta)$ (Average)
(d) $3cos(x^{\top}\theta)$ (Weak)
(e) $5cos(x^{\top}\theta)$ (Average)
(f) $5cos(x^{\top}\theta)$ (Weak)
(g) $10cos(x^{\top}\theta)$ (Average)
(h) $10cos(x^{\top}\theta)$ (Weak)
Figure 6: Comparisons of cumulative regret (average and weak) of different dueling bandits algorithms for different form of Cosine function, i.e., $cos(3x^{\top}\theta)$ and $bcos(x^{\top}\theta)$, where $b=\{3,5,10\}$.
C.1 Computational Efficiency and Scalability
To discuss the computational efficiency and scalability of our proposed algorithms, we consider the following two key aspects:
Size of the neural network (NN):  The primary computational complexity of our proposed algorithms from the NN used to estimate the latent reward function. Let $d$ be the dimension of the context-arm feature vector (input dimension for the NN). Assume the NN has $L$ hidden layers, each with $m$ neurons. Then, the inference cost from the input layer to the first hidden layer is $O(dm)$, the inference cost for hidden layers is $O(Lm^{2})$, and the inference cost for the final layer is $O(m)$. Therefore, the overall inference cost for each context-arm pair is $O(dm+Lm^{2}+m)$. Note that $p=dm+Lm^{2}+m$ is the total number of NN parameters. The training time for the NN is $O(ECLm^{2})$, where $E$ is the number of training epochs and $C$ is the number of context-arm pairs observations. Therefore, choosing the appropriate size of NN for the given problem is very important, as having smaller NNs may not accurately approximate the latent reward function, while larger NNs can incur significant training and inference costs.
Number of contexts and arms:  Let $K$ be the number of arms and $p$ be the total number of NN parameters. Since we use the gradients with respect to the current estimate of NN as context-arm features, the gradient computation cost for each context is $O(Kdp)$, where $d$ is the dimension of the context-arm feature vector. The computational costs for computing reward estimate and confidence terms are $O(Kp)$ and $O(Kp^{2})$, respectively. For selection arms, the computational cost for the first arm is $O(Kp+K)$, which includes estimating the reward for each context-arm pair $O(Kp)$ and then selecting the arm with the highest estimated reward $O(K)$; and the computational cost for the second arm is $O(Kp+(K-1)p^{2})$, i.e., computing estimated reward $O(Kp)$ and confidence terms $O((K-1)p^{2})$ for all arm with respect to the first selected arm Thus, the overall computational cost of our proposed algorithms for selecting a pair of arms for each context is $O(Kdp+Kp^{2})$. Since each context-arm pair is independent of others, each arm’s gradients, reward estimate, and optimistic term can be computed in parallel. Consequently, the computational cost for selecting a pair of arms for each context can reduced to $O(dp+p^{2})$.
C.2 Comparision with existing contextual dueling bandit algorithms
Saha (2021) proposed a contextual linear dueling bandits algorithm using a logistic link function with pairwise and subset-wise preference feedback. Whereas Bengs et al. (2022) generalized the setting to the contextual linear stochastic transitivity model, and Li et al. (2024) proposed an algorithm based on Feel-Good Thompson Sampling (Zhang, 2022). These existing works only consider the linear reward function, which may not be practical in many real-life applications. Our work fills this gap in the literature and generalizes the existing setting by considering the non-linear reward function in contextual dueling bandits. Furthermore, the arm-selection strategies are different, as Saha (2021) and Li et al. (2024) select the most informative pair of arms, and Bengs et al. (2022) select the first arm optimistically and select the second arm that beats the first arm optimistically. In contrast, we select the first arm greedily to ensure the best-performing arm is always selected and then the second arm optimistically to focus on exploration, making the arm-selection process computationally efficient as our algorithms only need to compare $K-1$ pairs instead of all possible pairs (i.e., $O(K^{2})$). Since we use an NN-based reward estimator, our optimistic term calculations for the context-arm pairs differ from the existing work. Therefore, our regret analysis uses techniques different from existing algorithms for contextual dueling bandits.
C.3 What if first and second arm are same
Since our proposed algorithms have a sub-linear regret guarantee, they will eventually select the best arm and recommend it in duels. This implies that for a given context, both the first and second arms chosen may ultimately be the same, representing the best-performing arm. Therefore, the first and second arms can be the same in our proposed algorithms. However, observations in which both arms are the same do not provide any useful preference information for estimating the latent reward function, as no comparison is being made. Including such observations for estimating latent reward function would only contribute to noise and result in a constant loss. Thus, we have to exclude these observations from the reward estimation.
Appendix D Neural Contextual Bandits with Binary Feedback
We extend our results to the neural contextual bandit problem in which a learner only observes binary feedback for the selected arms (note that the learner only selects one arm in every iteration). Observing binary feedback is very common in many real-life applications, e.g., click or not in online recommendation and treatment working or not in clinical trials (Li et al., 2017; Faury et al., 2020).
Contextual bandits with binary feedback.
Performance measure.
D.1 Reward function estimation using neural network and our algorithms
UCB-based algorithm.
1:  Tuning parameters: $\delta\in(0,1)$ and $\lambda>0$
2:  for $t=1,\ldots,T$ do
6:     Observe preference feedback binary $y_{t}$
7:  end for
TS-based algorithm.
Regret analysis.
We use the above fact to prove the following confidence ellipsoid result as done in linear reward function (Li et al., 2017; Jun et al., 2017; Faury et al., 2020).
Theorem 4.
for all $x\in\mathcal{X}_{t},t\in[T]$.
Theorem 5 (NCBF-UCB).
$\mathfrak{R}_{T}=\widetilde{O}\left(\left(\frac{\sqrt{\widetilde{d}_{b}}}{\kappa_{\mu}}+B\sqrt{\frac{\lambda}{\kappa_{\mu}}}\right)\sqrt{T\widetilde{d}_{b}}\right)$
Theorem 6 (NCBF-TS).
Note that in terms of asymptotic dependencies (ignoring the log factors), our UCB- and TS- algorithms have similar growth rates. All missing proofs and additional details are in Section D.2.
Comparison with Neural Bandits. The regret upper bounds of our NCBF-UCB and NCBF-TS algorithms are worse than the regret of NeuralUCB Zhou et al. (2020) and NeuralTS Zhang et al. (2021): $\widetilde{O}(\widetilde{d}_{b}\sqrt{T})$ (with $\kappa_{\mu}=1$) because of our extra dependency on $\kappa_{\mu}$ and $L_{\mu}$. [^3] Note that our effective dimension $\widetilde{d}_{b}$ is defined using $\mathbf{H}_{b}$ Eq. 36, while the effective dimension $\widetilde{d}^{\prime}$ in Zhou et al. (2020) and Zhang et al. (2021) are defined using $\mathbf{H}$. However, as we have discussed in Footnote 2, $\widetilde{d}^{\prime}$ has the same order of growth as $\log\det\left(\mathbf{H}_{b}/\lambda+\mathbf{I}\right)$. So, our regret upper bounds are comparable with those from Zhou et al. (2020) and Zhang et al. (2021). Specifically, note that $\kappa_{\mu}<1$, therefore, the regret bounds in 5 and 6 are increased as a result of the dependency on $\kappa_{\mu}$. In addition, the dependency on $L_{\mu}$ also places an extra requirement on the width $m$ of the NN. Therefore, our regret bounds are worse than that of standard neural bandit algorithms that do not depend on $\kappa_{\mu}$ and $L_{\mu}$. This can be attributed to the additional difficulty of our problem setting, i.e., we only have access to binary feedback, whereas standard neural bandits Zhou et al. (2020); Zhang et al. (2021) can use continuous observations.
D.2 Theoretical Analysis